# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 2.1 Saddle point problem formulations
Denoting $\Pi:\mathbb{R}^{md}\to\mathbb{R}^{md}$ as the projection such that for any $X\in\mathbb{R}^{md}$, $\Pi(X)_{v}=\frac{1}{m}\sum_{v^{\prime}\in V}x_{v^{\prime}}$, we have $KX=\mathbf{0}$ if and only if $\Pi X=X$. Thus, Assumption 2.2 implies that
$K^{*}(KK^{*})^{\dagger}K=I-\Pi.$
For convenience, we further decompose $K_{s}X=\sum_{v\in V}K_{s,v}x_{v}$ for $K_{s,v}:\mathbb{R}^{d}\to\mathbb{R}^{n_{s}}$.
Penalty. In ($\mathcal{P}_{r}$), $R_{s}:\mathbb{R}^{n_{s}}\to\overline{\mathbb{R}}$ is a regularization term, penalizing the deviation of $K_{s}X$ from $\mathbf{0}$. We further define $R:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ as $R(y_{1},\ldots,y_{S})=\sum_{s=1}^{S}R_{s}(y_{s})$. Recall that the Fenchel conjugate of $R_{s}$ is defined as
$R_{s}^{*}(y_{s})=\sup_{y_{s}^{\prime}\in\mathbb{R}^{n_{s}}}\langle y_{s},y_{s}^{\prime}\rangle-R_{s}(y^{\prime}_{s}),$
and it is easy to see $R^{*}(Y)=\sum_{s=1}^{S}R^{*}_{s}(y_{s})$ for $Y=(y_{s})_{s\in[S]}$. We make the following assumption.
Assumption 2.3.
For each $s\in[S]$, $R_{s}:\mathbb{R}^{n_{s}}\to\overline{\mathbb{R}}$,
  1. 1.
    $R_{s}$ is proper, convex, and lower-semicontinuous;
  2. 2.
    $R_{s}(y_{s})\geq 0$ for all $y_{s}\in\mathbb{R}^{n_{s}}$, and $R_{s}(y_{s})=0$ if and only if $y_{s}=\mathbf{0}$.
@@
As an example, if for all $s\in[S]$, $R_{s}$ is the characteristic function of the set $\{\mathbf{0}\}$, i.e. $R_{s}(\mathbf{0})=0$ and $R_{s}(y_{s})=\infty$ for $y_{s}\neq\mathbf{0}$, then ($\mathcal{P}_{r}$) is equivalent to ($\mathcal{P}$), and $R_{s}^{*}(y_{s})=0$ for all $y_{s}\in\mathbb{R}^{n_{s}}$. As another example, $R_{s}$ can be any scaled norm, for instance $R_{s}(y_{s})=\lambda\|y_{s}\|_{p}$ for some $p\geq 1$ and $\lambda>0$, then $R_{s}^{*}(y_{s})=0$ for $\|y_{s}\|_{q}\leq\lambda$ and $R_{s}^{*}(y_{s})=\infty$ otherwise, where $\|\cdot\|_{q}$ is the dual norm of $\|\cdot\|_{p}$ (i.e. $p^{-1}+q^{-1}=1$). That is $R_{s}^{*}$ is the characteristic function of the dual-norm-ball of size $\lambda$. We would like to point out that the first condition in Assumption 2.3 is standard in the PDHG literature [chambolle_ergodic_2016, chambolle_first-order_2011]. We discuss further requirements on the choices of $R_{s}$ in Section 2.3.
Under the first part of Assumption 2.3, we have $R_{s}=R_{s}^{**}$ (Theorem 11.1 [rockafellarVariational]). Thus, ($\mathcal{P}_{r}$) can be equivalently formulated as the following saddle point problem:
$\min_{X\in\overline{\mathcal{X}}}\max_{Y\in\mathbb{R}^{n}}\sum_{s=1}^{S}\langle K_{s}X,y_{s}\rangle+F(X)-\sum_{s=1}^{S}R_{s}^{*}(y_{s}).$ ($\mathcal{P}_{s}$)
@@
Performance measure. To measure the performance of $X\in\overline{\mathcal{X}}$, following [lan_communication-efficient_2020], we consider the $(\epsilon,\delta)$-solution, satisfying the following conditions
@@
That is, $X$ is $\epsilon$-suboptimal in terms of the objective value, while violating the consensus constraints by at most $\delta$ [^2] In [lan_communication-efficient_2020], $\|KX\|\leq\delta$ is used instead of $\|(I-\Pi)X\|$, and $K$ is assumed to be the Laplacian matrix for the underlying graph of communication. However, we use a generic $K$ satisfying condition 2.2. In particular, for any $K$ that is a valid choice, $\lambda K$ is also valid for any $\lambda\neq 0$. Thus, it makes sense to “normalize” $K$, and we use $I-\Pi=K^{*}(KK^{*})^{\dagger}K$.
To solve ($\mathcal{P}$), we resort to the primal dual formulation ($\mathcal{P}_{s}$), where the common measure of performance of $(X,Y)$ is the duality gap, defined as $G:\mathcal{Z}\times\mathcal{Z}\to\overline{\mathbb{R}}$ where $\mathcal{Z}=\overline{\mathcal{X}}\times\mathbb{R}^{n}$ as
$G(X,Y;X^{\prime},Y^{\prime}):=$ $\langle KX,Y^{\prime}\rangle+F(X)-R^{*}(Y^{\prime})$
$-\left\{\langle KX^{\prime},Y\rangle+F(X^{\prime})-R^{*}(Y)\right\}.$ (3)
Our algorithms are guaranteed to find a pair $Z=(X,Y)$ such that $\sup_{Z^{\prime}\in\mathcal{Z}}G(Z;Z^{\prime})$ is small (Theorems 3.1 and 4.1). To transfer such duality gap guarantee back to $(\epsilon,\delta)$-solution guarantee, in Section 2.3, we propose additional requirements for the regularization $R$.
[…]
Assumption 2.4.
$\min_{y_{s}\in\mathsf{dom}(R^{*}_{s})}R_{s}^{*}(y_{s})+\langle g,y_{s}\rangle+D_{w^{y}_{s}}(y_{s},\overline{y}_{s}).$
For any $g\in\mathbb{R}^{d}$, the following problem can be solved exactly:
$\min_{x\in\mathcal{X}}\langle g,x\rangle+w^{x}(x).$
@@ 2.2 Agents, communication, and additional requirements on $K$
We assume that for any pair $(s,v)\in[S]\times V$ such that $K_{s,v}\neq\mathbf{0}$, $\mathsf{Agent}(x_{v})$ and $\mathsf{Agent}(y_{s})$ can communicate (in both directions). For instance, all agents might be nodes in a connected graph with vertices $[S]\cup V$ (representing $S$ dual agents and $m$ primal agents), and communication can be realized through edges (directly) or through paths (i.e. with the help of intermediate agents). In particular, since the graph is connected, any pair can communicate, but the resources consumed and/or time taken by communication between different pairs could be (significantly) different.
At this point, we abstract away from how such communication is realized, and leave the discussion of the costs of communication to Sections 3.5.1 and 3.5.2. Below, we provide two such realizations: decentralized and hierarchical, and provide examples in Figure 1.
@@
As an example, let $W\in\mathbb{R}^{V\times V}$ be a doubly stochastic matrix such that $W_{v,v^{\prime}}\neq 0$ only if $\{v,v^{\prime}\}\in E$ or $v=v^{\prime}$, and $\ker(I-W)=\mathsf{Span}(\mathbf{1})$ (and so $K:=(I-W)\otimes I_{d}$ satisfies Assumption 2.2). We can choose $S=m$, $n_{s}=d$, and decompose $K$ as $K_{s}:=(I-W)_{s}\otimes I_{d}$,
@@
Hierarchical setting. In this setup, we assume that there is an underlying tree with nodes $[S]\cup V$, where all non-leaf nodes ($[S]$) correspond to dual agents and all leaf nodes ($V$) correspond to primal agents. Each non-leaf node can communicate with its child nodes directly. Precisely, for $s\in[S]$, we use $\mathsf{Chi}(s)\subset[S]\cup V$ to denote the child nodes of $\mathsf{Agent}(y_{s})$, and $\mathsf{Des}(s)\subset V$ to denote all primal agents in the subtree rooted at $\mathsf{Agent}(y_{s})$.
For convenience, for each $s\in[S]$, we denote the “mean” of all descendants of $\mathsf{Agent}(y_{s})$ as $\overline{x}_{s}=|\mathsf{Des}(s)|^{-1}\sum_{j\in\mathsf{Des}(s)}x_{j}$. Then, consider $K_{s}:\mathbb{R}^{md}\to\mathbb{R}^{|\mathsf{Chi}(s)|d}$ defined as
@@
Then, it is easy to see that $K$ satisfies Assumption 2.2, and since $\overline{x}_{j}$ can be computed in a bottom up manner, $\{K_{s}\}_{s\in[S]}$ can be realized through this tree. In addition, the set of $\{K_{s}\}_{s\in[S]}$ admits the following orthogonality properties which will be useful in choosing $R$. We defer the proof to Appendix A.
@@
Let $\{K_{s}\}_{s\in[S]}$ be as defined in (4). Then for $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$. In addition, denoting $\Pi_{s}:=K_{s}^{*}(K_{s}K_{s}^{*})^{\dagger}K_{s}$, we have for any $\widetilde{X},\widehat{X}\in\mathbb{R}^{md}$
@@
2.3 Requirements for $R$
Recall that when $R$ is the characteristic function of $\{\mathbf{0}\}$, the penalized formulation ($\mathcal{P}_{r}$) and its primal-dual version ($\mathcal{P}_{s}$) are equivalent to ($\mathcal{P}$). In this section, we discuss the requirements for $R$ such that the duality gap provides upper bounds on the suboptimality of the objective value and the violation of the consensus constraints.
@@ 2.3 Requirements for $R$
Under Assumption 2.3, we have for any $\widehat{X}\in\overline{\mathcal{X}}$ such that $K\widehat{X}=\mathbf{0}$,
@@
In particular, under Assumption (2.1), if $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$, then $F(X)\leq F(X^{*})+\epsilon$.
@@
Notice that by Assumption (2.3),
@@
$\langle K\widehat{X},Y\rangle+F(\widehat{X})-R^{*}(Y)\leq F(\widehat{X}),\quad\forall~{}\widehat{X}\in\overline{\mathcal{X}},~{}K\widehat{X}=\mathbf{0}.$
@@
For convenience, we denote $\sigma_{\min}^{+}(K_{s})=\min_{X\in\mathbb{R}^{md},~{}\Pi_{s}X\neq\mathbf{0}}\frac{\|K_{s}X\|_{*}}{\|\Pi_{s}X\|}$, where the numerator uses the dual norm to the norm in $\mathbb{R}^{n_{s}}$ and the denominator uses the norm in $\mathbb{R}^{md}$. As an example, when all norms are $l_{2}$ norms, $\sigma_{\min}^{+}(K_{s})$ is the smallest non-zero singular value of $K_{s}$.
@@ 2.3.1 Requirements on $R$ under orthogonaltiy
Under Assumption (2.3), further assuming that for any $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$, and for each $s\in[S]$, denoting
$\Pi_{s}=K_{s}^{*}(K_{s}K_{s}^{*})^{\dagger}K_{s},\quad a_{s}\geq\sup_{X^{\prime}\in\overline{\mathcal{X}},~{}KX^{\prime}=\mathbf{0}}\|\Pi_{s}\nabla F(X^{\prime})\|_{*},$
where $\nabla F:\overline{\mathcal{X}}\to\mathbb{R}^{md}$ is an arbitrary subgradient oracle, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v})$. If Assumption 2.1 also holds and $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$,
@@
In Lemma 2.3 (Corollary 2.1 below), the superscript $ccv$ means $\{R^{ccv}_{s}\}_{s\in[S]}$ ($\{\widehat{R}^{ccv}_{s}\}_{s\in[S]}$) are designed to provide guarantees on the consensus constraint violation, and the superscript $prj$ means $\{R^{prj}_{s}\}_{s\in[S]}$ ($\{\widehat{R}^{prj}_{s}\}_{s\in[S]}$) are designed to provide guarantees the projected solution $\Pi X$.
@@
We would like to point out that in (2.3.1), $\langle\nabla F(\widetilde{X}),\Pi_{s}X\rangle$ is upper bounded using $\|\nabla F(\widetilde{X})\|_{*}\cdot\|\Pi_{s}X\|$. A tighter upper bound could be obtained if one has more information about the set $\mathcal{G}_{s}:=\{\Pi_{s}\nabla F(X^{\prime}),~{}X^{\prime}\in\overline{\mathcal{X}},~{}KX^{\prime}=\mathbf{0}\}$. Indeed, $\langle\nabla F(\widetilde{X}),\Pi_{s}X\rangle\leq\sup_{G_{s}\in\mathcal{G}_{s}}\langle G_{s},\Pi_{s}X\rangle$, and so the inner product can be bounded using the support function of the set $\mathcal{G}_{s}$.

### v2

@@ 4.1.1 Consensus constraints
As an example, denoting $\Pi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as the projection such that for any $X\in\mathbb{R}^{d}$, $\Pi(X)_{v}=\frac{1}{m}\sum_{v^{\prime}\in V}x_{v^{\prime}}$, then we can take $K=I-\Pi$. Moreover, for any $K$ satisfying Assumption 4.1, $K^{*}(KK^{*})^{\dagger}K=I-\Pi$ holds.
We point out that the kernel condition in Assumption 4.1 is the only requirement we impose on $K$. However, to make sure the resulting primal-dual algorithm can be implemented in a distributed fashion, additional sparsity requirements are needed depending on how agents communicate with each other. We provide details, examples, and the rationale behind the block-decomposable formulation in Section 4.2.
4.1.2 Penalties
In ($\mathcal{P}_{d}^{lift}$), we assume that $R_{s}:\mathbb{R}^{n_{s}}\to\overline{\mathbb{R}}$ is proper, convex, and lower-semicontinous, and $R_{s}(\mathbf{0})=0$. Under these conditions, ($\mathcal{P}_{b}$) (and thus ($\mathcal{P}_{d}^{lift}$)) admits a primal-only formulation ($\mathcal{P}_{p}$). In this formulation, $R_{s}(K_{s}X)$ becomes a penalty term, penalizing the deviation of $K_{s}X$ from $\mathbf{0}$. We further define $R:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ as $R(y_{1},\ldots,y_{S})=\sum_{s=1}^{S}R_{s}(y_{s})$.
@@ 4.1.2 Penalties
As an example, if for all $s\in[S]$, $R_{s}$ is the characteristic function of the set $\{\mathbf{0}\}$, i.e. $R_{s}(\mathbf{0})=0$ and $R_{s}(y_{s})=\infty$ for $y_{s}\neq\mathbf{0}$, then ($\mathcal{P}_{d}^{lift}$) is equivalent to ($\mathcal{P}_{d}$), and $R_{s}^{*}(y_{s})=0$ for all $y_{s}\in\mathbb{R}^{n_{s}}$. As another example, $R_{s}$ can be any scaled norm, for instance $R_{s}(y_{s})=\lambda\|y_{s}\|_{p}$ for some $p\geq 1$ and $\lambda>0$, then $R_{s}^{*}(y_{s})=0$ for $\|y_{s}\|_{q}\leq\lambda$ and $R_{s}^{*}(y_{s})=\infty$ otherwise, where $\|\cdot\|_{q}$ is the dual norm of $\|\cdot\|_{p}$ (i.e. $p^{-1}+q^{-1}=1$). That is $R_{s}^{*}$ is the characteristic function of the dual-norm-ball of size $\lambda$.
Notice that with penalties different from the characteristic functions of $\{\mathbf{0}\}$, the problem ($\mathcal{P}_{d}^{lift}$) becomes a “relaxation” of ($\mathcal{P}_{d}$), thereby are not equivalent. As one can imagine, the two formulations get closer as the consensus constraints are penalized more. In fact, there are two tensions in choosing good penalties. Take $S=1$ and $R(Y)=\lambda\|Y\|$ as an example.
  • On one hand, with larger penalty on the consensus constraint violation (i.e. larger $\lambda$), ($\mathcal{P}_{d}^{lift}$) will become a better proxy for ($\mathcal{P}_{d}$); this encourages larger $\lambda$.
  • On the other hand, the diameter of $\mathsf{dom}(R^{*})$ is $O(\lambda)$, and as suggested by Corollaries 3.3 and 3.4, the complexities of our algorithms are $O(\lambda)$; this encourages smaller $\lambda$.
In Section 4.3, we propose a set of conditions on the penalties to achieve a balance between these two tensions, such that “good solutions” to ($\mathcal{P}_{d}^{lift}$) and ($\mathcal{P}_{b}$) (as measured using the duality gap (2)) are also “good solutions” to ($\mathcal{P}_{d}$) (as measured by objective value suboptimality and constraint violation in (18)), and the resulting algorithms have favorable dependence on the problem parameters.
4.1.3 Performance measure
@@
To measure the performance of $X\in{\mathcal{X}}$, following [27], we consider the $(\epsilon,\delta)$-solution, satisfying the following conditions
@@
That is, $X$ is $\epsilon$-suboptimal in terms of the objective value, while violating the consensus constraints by at most $\delta$. [^5] In [27], $\|KX\|\leq\delta$ is used instead of $\|(I-\Pi)X\|$, and $K$ is assumed to be the Laplacian matrix for the underlying graph of communication. However, we use a generic $K$ satisfying condition 4.1. In particular, for any $K$ that is a valid choice, $\lambda K$ is also valid for any $\lambda\neq 0$. Thus, it makes sense to “normalize” $K$, and we use $I-\Pi=K^{*}(KK^{*})^{\dagger}K$. Nevertheless, our algorithms have performance guarantees on the duality gap of the saddle point fromulation. To transfer such duality gap guarantee back to $(\epsilon,\delta)$-solution guarantee, in Section 4.3, we propose additional requirements for the regularization $R$.
@@ 4.2 Agents, communication, and additional requirements on $K$
We assume that for any pair $(s,v)\in[S]\times V$ such that $K_{s,v}\neq\mathbf{0}$, $\mathsf{Agent}(x_{v})$ and $\mathsf{Agent}(y_{s})$ can communicate (in both directions). For instance, all agents might be nodes in a connected graph with vertices $S]\cup V$ (representing $S$ dual agents and $m$ primal agents), and communication can be realized through edges (directly) or through paths (i.e. with the help of intermediate agents). In particular, since the graph is connected, any pair can communicate, but the resources consumed and/or time taken by communication between different pairs could be (significantly) different.
At this point, we abstract away from how such communication is realized, and leave the discussion of the costs of communication to Section 4.4. Below, we provide two such realizations: decentralized and hierarchical, and provide examples in Figure 2.
@@
As an example, let $W\in\mathbb{R}^{V\times V}$ be a doubly stochastic matrix such that $W_{v,v^{\prime}}\neq 0$ only if $\{v,v^{\prime}\}\in E$ or $v=v^{\prime}$, and $\ker(I-W)=\mathsf{Span}(\mathbf{1})$ (and so $K:=(I-W)\otimes I_{\overline{d}}$ satisfies Assumption 4.1). We can choose $S=m$, $n_{s}=d$, and decompose $K$ as $K_{s}:=(I-W)_{s}\otimes I_{\overline{d}}$,
@@
Hierarchical setting. In this setup, we assume that there is an underlying tree with nodes $S]\cup V$, where all non-leaf nodes ($[S]$) correspond to dual agents and all leaf nodes ($V$) correspond to primal agents. Each non-leaf node can communicate with its child nodes directly. Precisely, for $s\in[S]$, we use $\mathsf{Chi}(s)\subset[S]\cup V$ to denote the child nodes of $\mathsf{Agent}(y_{s})$, and $\mathsf{Des}(s)\subset V$ to denote all primal agents in the subtree rooted at $\mathsf{Agent}(y_{s})$.
For convenience, for each $s\in[S]$, we denote the “mean” of all descendants of $\mathsf{Agent}(y_{s})$ as $\overline{x}_{s}=|\mathsf{Des}(s)|^{-1}\sum_{j\in\mathsf{Des}(s)}x_{j}$. Then, consider $K_{s}:\mathbb{R}^{d}\to\mathbb{R}^{|\mathsf{Chi}(s)|{\overline{d}}}$ defined as
@@
Then, it is easy to see that $K$ satisfies Assumption 4.1, and since $\overline{x}_{j}$ can be computed in a bottom up manner, $\{K_{s}\}_{s\in[S]}$ can be realized through this tree. In addition, the set of $\{K_{s}\}_{s\in[S]}$ admits the following orthogonality properties which will be useful in choosing $R$. We defer the proof to Appendix B.
@@
Let $\{K_{s}\}_{s\in[S]}$ be as defined in (19). Then for $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$. In addition, denoting $\Pi_{s}:=K_{s}^{*}(K_{s}K_{s}^{*})^{\dagger}K_{s}$, we have for any $\widetilde{X},\widehat{X}\in\mathbb{R}^{d}$
@@
4.3 Requirements for $R$ and function similarity
Recall that when $R$ is the characteristic function of $\{\mathbf{0}\}$, the penalized formulation ($\mathcal{P}_{d}^{lift}$) is equivalent to ($\mathcal{P}_{d}$). In this section, we discuss the requirements for $R$ such that the duality gap provides upper bounds on the suboptimality of the objective value and the violation of the consensus constraints.
First, we have the following upper bounds on the suboptimality of the objective value.
@@ 4.3 Requirements for $R$ and function similarity
For any $\widehat{X}\in{\mathcal{X}}$ such that $K\widehat{X}=\mathbf{0}$,
@@
In particular, if $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$, then $F(X)\leq F(X^{*})+\epsilon$, where $X^{*}=(x^{*})_{v\in V}$ and $x^{*}$ is an optimal solution to ($\mathcal{P}_{d}$).
@@
Recall that we have
@@
$\langle K\widehat{X},Y\rangle+F(\widehat{X})-R^{*}(Y)\leq F(\widehat{X}),\quad\forall~\widehat{X}\in{\mathcal{X}},~K\widehat{X}=\mathbf{0}.$
@@
For convenience, we denote $\sigma_{\min}^{+}(K_{s})=\min_{X\in\mathbb{R}^{d},~\Pi_{s}X\neq\mathbf{0}}\frac{\|K_{s}X\|_{*}}{\|\Pi_{s}X\|}$, where the numerator uses the dual norm to the norm in $\mathbb{R}^{n_{s}}$ and the denominator uses the norm in $\mathbb{R}^{d}$. As an example, when all norms are $l_{2}$ norms, $\sigma_{\min}^{+}(K_{s})$ is the smallest non-zero singular value of $K_{s}$.
@@ 4.3.1 Requirements on $R$ under orthogonaltiy
Further assume that for any $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$, and for each $s\in[S]$, denoting
$\Pi_{s}=K_{s}^{*}(K_{s}K_{s}^{*})^{\dagger}K_{s},\quad a_{s}\geq\sup_{X^{\prime}\in{\mathcal{X}},~KX^{\prime}=\mathbf{0}}\|\Pi_{s}\nabla F(X^{\prime})\|_{*},$
where $\nabla F:{\mathcal{X}}\to\mathbb{R}^{d}$ is an arbitrary subgradient oracle, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v})$. If $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$,
@@
In Lemma 4.3 (and Corollary 4.1 below), the superscript ccvccv means $\{R^{ccv}_{s}\}_{s\in[S]}$ ($\{\widehat{R}^{ccv}_{s}\}_{s\in[S]}$) are designed to provide guarantees on the consensus constraint violation, and the superscript prjprj means $\{R^{prj}_{s}\}_{s\in[S]}$ ($\{\widehat{R}^{prj}_{s}\}_{s\in[S]}$) are designed to provide guarantees the projected solution $\Pi X$.
@@
We would like to point out that in (4.3.1), $\langle\nabla F(\widetilde{X}),\Pi_{s}X\rangle$ is upper bounded using $\|\nabla F(\widetilde{X})\|_{*}\cdot\|\Pi_{s}X\|$. A tighter upper bound could be obtained if one has more information about the set $\mathcal{G}_{s}:=\{\Pi_{s}\nabla F(X^{\prime}),~X^{\prime}\in{\mathcal{X}},~KX^{\prime}=\mathbf{0}\}$. Indeed, $\langle\nabla F(\widetilde{X}),\Pi_{s}X\rangle\leq\sup_{G_{s}\in\mathcal{G}_{s}}\langle G_{s},\Pi_{s}X\rangle$, and so the inner product can be bounded using the support function of the set $\mathcal{G}_{s}$.